how to solve this2. You decide to buy an ice cream cart to sell frozen desserts on Friday afternoons outside the Questrom building. The desser
Answer #1

NPV = present value of future cash flows - initial cash outflow

cash flow per week = number of units sold x (selling price - cost)

= 300 x (5 - 3)

= $600

cash flows are even through out the entire year

given discount rate = 4%

the above rate is per year

weekly rate = 4 / 52 = 0.076923% (number of periods n = 52)

present value of future cash flows = 600 x PVIFA( r = 0.076923% ; n = 52)

= 600 x 50.954523

= 30,573 (rounded to nearest dollar)

so NPV = 30573 - 10000

= $20,573